OpenStudy (mallorysipp234):

Arthur is testing the effectiveness of a new acne medication. There are one hundred people with acne in the study. Forty patients received the acne medication and sixty other patients did not receive treatment. Fifteen of the patients who received the medication reported clearer skin at the end of the study. Twenty of the patients who did not receive medication reported clearer skin at the end of the study. What is the probability that a patient chosen at random from this study took the medication, given that they reported clearer skin? 15% 33% 38% 43%

OpenStudy (mallorysipp234):

@thomaster @sososmart @wolf1728

OpenStudy (mallorysipp234):

@ganeshie8

OpenStudy (mallorysipp234):

@IAmSinged

OpenStudy (mallorysipp234):

@ranga please help

OpenStudy (anonymous):

We know that 15 people took the medication had clearer skin. We know that 20 people didn't take the medication and had clearer skin. So the number of people who had clearer skin is 35. The probably is: (people on medication AND clearer skin)/(people with clearer skin) = 15/35 = 43%. 4th option.